<!DOCTYPE HTML PUBLIC "-//W3C//DTD HTML 4.01 Transitional//EN">
<html>
  <head>

    <meta http-equiv="content-type" content="text/html; charset=ISO-8859-1">
  </head>
  <body>
    <title>BTeam-2011-01-12</title>
    <meta http-equiv="Content-Type" content="text/html;
      charset=ISO-8859-1">
    <meta name="generator" content="Org-mode">
    <meta name="generated" content="2011/01/12 12:50:38">
    <meta name="author" content="Todd Satogata">
    <style type="text/css">
  html {
        font-family: Times, serif;
        font-size: 12pt;
  }
  .title { text-align: center; }
  .todo  { color: red; }
  .done { color: green; }
  .timestamp { color: grey }
  .timestamp-kwd { color: CadetBlue }
  .tag { background-color:lightblue; font-weight:normal }
  .target { background-color: lavender; }
  pre {
        border: 1pt solid #AEBDCC;
        background-color: #F3F5F7;
        padding: 5pt;
        font-family: courier, monospace;
  }
  table { border-collapse: collapse; }
  td, th {
        vertical-align: top;
        <!--border: 1pt solid #ADB9CC;-->
  }
</style>BTeam -<br>
    <br>
    Here is the BTeam agenda for today, and a copy of last week's
    minutes. We should plan to prioritize modest study time for tomorrow
    and next Tuesday depending on study time availability since high
    current setup took a bite out of ABUs yesterday.<br>
    <br>
    -Todd<br>
    <h1 class="title">BTeam-2011-01-12</h1>
    <div class="outline-2">
      <h2 id="sec-1">1 2011-01-12 Wed 14:00 BTeam Meeting Agenda</h2>
      <div class="outline-3">
        <h3 id="sec-2">1.1 Optics-on-Call Report: Jay</h3>
        <p>A long week returning to hall C high current.
        </p>
      </div>
      <div class="outline-3">
        <h3 id="sec-3">1.2 Optics-on-Call schedule: Todd</h3>
        The usual <a
href="http://opsweb.acc.jlab.org/CSUEApps/weboncall/groupSchedule.html?Optics">website
          is here</a> for Web On Call for Optics. I plan to
        keep this 7-team rotation (changing at BTeam meeting) until the
        summer down, though we may add an extra team to bring Vasily
        Morozov on board. Teams may also swap coverage with consensus of
        both teams.<br>
        <ul>
          <li>Jan 12-19 (USPAS) Mike Spata/Ryui</li>
          <li>Jan 19-26 (USPAS) Yuhong/Tief</li>
          <li>Jan 26-Feb 2 Byung/Yves</li>
          <li>Feb 2-Feb 9 AlexB/Balsa (or Vasily)</li>
          <li>Feb 9-Feb 16 Geoff/Shahid</li>
          <li>Feb 16-Feb 23 Alicia/Reza</li>
          <li>Feb 23-Mar 2 Jay/Todd
          </li>
        </ul>
      </div>
      <div class="outline-3">
        <h3 id="sec-4">1.3 Beam Studies Report</h3>
        <div class="outline-4">
          <h4 id="sec-5">1.3.1 Mott Measurements: Grames/Suleiman</h4>
          <p><a href="http://cebaf.jlab.org/atlis/task/10588">ATLis
              10589</a>: Completed characterization of new detectors
            after
            shutdown installation, spin rotator checkout.
          </p>
        </div>
        <div class="outline-4">
          <h4 id="sec-6">1.3.2 Qweak Injector Studies/Parity Quality
            Check: Dalton</h4>
          <p><a href="https://cebaf.jlab.org/atlis/task/10625">ATLis
              10625</a>: Shifted to Thursday owl 2010-01-06 at Mark's
            request.
            But also should be revisited at high current.
          </p>
        </div>
        <div class="outline-4">
          <h4 id="sec-7">1.3.3 Backphasing Bunch Length Measurement
            @100kV: Hofler</h4>
          <p><a href="https://cebaf.jlab.org/atlis/task/10684">ATLis
              10684</a>: Abbreviated version performed 2011-01-11 as
            part of
            high current restoration. Modifications made to test plan
            for
            future testing. Should be measured again before we leave 100
            kV
            injector voltage.
          </p>
        </div>
        <div class="outline-4">
          <h4 id="sec-8">1.3.4 Pathlength vs MOmod: Benesch</h4>
          <p>Studied during day 2010-01-05 during setup. NL response
            seems to
            be about 80% of SL. Haven't looked at owl MOmod. Looks like
            20 mV/deg
            in NL, 25 mV/deg in SL, different than last measurement.
            Will continue
            to evaluate. Further comment?
          </p>
        </div>
      </div>
      <div class="outline-3">
        <h3 id="sec-9">1.4 Beam Studies Planning</h3>
        <div class="outline-4">
          <h4 id="sec-10">1.4.1 LOCO Model Calibration Tests in ARC3:
            Roblin</h4>
          <p><a href="http://cebaf.jlab.org/atlis/task/10700">ATLis
              10700</a>: Yves has compared some of his measurements in
            ARC9
            from the 2010-12-23 <a
              href="https://cebaf.jlab.org/atlis/task/10511">ATLis 10511</a>
            to current and historical MQA
            curves; they show some discrepancies worth investigating
            with
            more measurements in other arcs. Another method is in
            development
            to handle non-FODO quad distributions in arcs 1 and 2.
          </p>
        </div>
        <div class="outline-4">
          <h4 id="sec-11">1.4.2 Determine multiharp masking requirements
            in tune/CW: Benesch</h4>
          <p><a href="https://cebaf.jlab.org/atlis/task/10647">ATLis
              10647</a>: New ATLis, not scheduled yet. Mainly hall C.
          </p>
        </div>
        <div class="outline-4">
          <h4 id="sec-12">1.4.3 Prebuncher control module swap, injector
            retune</h4>
          <p><a href="https://cebaf.jlab.org/atlis/task/10508">ATLis
              10508</a>: To reduce 9.5 kHz noise when running prebuncher
            above 2. This ATLis does not seem strongly motivated until
            we
            go back to high injector voltage.
          </p>
          <ul>
            <li><span class="todo">TODO</span> Should investigate now
              we're at 100 kV w/lower prebuncher setting.<br>
            </li>
          </ul>
        </div>
      </div>
      <div class="outline-3">
        <h3 id="sec-13">1.5 Arc Diagnostics: Benesch</h3>
        <div class="outline-4">
          <h4 id="sec-14">1.5.1 <span class="todo">TODO</span> Analyze
            raw harp profile data and write up recommendations.</h4>
        </div>
      </div>
      <div class="outline-3">
        <h3 id="sec-15">1.6 12GeV Pathlength Requirements Status:
          Tiefenback</h3>
        <div class="outline-4">
          <h4 id="sec-16">1.6.1 <span class="todo">TODO</span> Need
            conclusions document.</h4>
        </div>
      </div>
      <div class="outline-3">
        <h3 id="sec-17">1.7 AOB?</h3>
        <p>
======================================================================
        </p>
      </div>
    </div>
    <div class="outline-2">
      <h2 id="sec-18">2 2011-01-05 Wed 14:00 BTeam Meeting Agenda</h2>
      <div class="outline-3">
        <h3 id="sec-19">2.1 Optics-on-Call Report: Benesch<br>
        </h3>
        <p>Some bumps, but went away after post-new-optics magnet cycle.
          To get round spots on sync light monitors, Yan had to tune up
          injector optics by hand, ~10% changes. Still need to look at
          fopt with injector excitation from earlier this AM; has
          injector
          excitation been reconnected?
        </p>
      </div>
      <div class="outline-3">
        <h3 id="sec-20">2.2 Beam Studies Report</h3>
        <div class="outline-4">
          <h4 id="sec-21">2.2.1 Arc9 Transport using LOCO Methods:
            Roblin</h4>
          <p><a href="https://cebaf.jlab.org/atlis/task/10511">ATLis
              10511</a>: Completed on 12/23. Draft results look
            promising,
            delivered in PDF and at CASA weekly 1/5. Yves presented some
            of
            his PDF slides at the BTeam meeting too. This approach
            measures
            many difference orbits from many correctors, and adjusts
            model
            quad/dipole gradient families and BPM gains to match
            measurement.
          </p>
          <ul>
            <li><span class="todo">TODO</span> Questions about whether
              we can cross-compare arc/linac BPMs,<br>
              check recombiner gradients, which do not seem to be in
              families.
              Yves will also work on algorithm for improved method
              similar to
              RayTrace. Plan for new measuements for gradient curves,
              compare
              to Jay's spreadsheet/CED. Estimate is ~1h/measurement/arc.
              Meet
              early next week.
            </li>
          </ul>
        </div>
        <div class="outline-4">
          <h4 id="sec-22">2.2.2 Mott Measurements and Detector Swap:
            Grames</h4>
          <p><a href="https://cebaf.jlab.org/atlis/task/10588">ATLis
              10588</a> and <a
              href="https://cebaf.jlab.org/atlis/task/10588">ATLis 10587</a>:
            Completed 12/23 day after shutdown
            of rest of facility except for NL using 100 kV PS.
          </p>
        </div>
        <div class="outline-4">
          <h4 id="sec-23">2.2.3 Qweak Injector Studies: Mark Dalton</h4>
          <p><a href="https://cebaf.jlab.org/atlis/task/10625">ATLis
              10625</a>: Shifted to Thursday owl 2010-01-06 at Mark's
            request.
          </p>
        </div>
        <div class="outline-4">
          <h4 id="sec-24">2.2.4 Arc 1/2 New Optics 2010-01-03: Arne</h4>
          <p><a href="https://cebaf.jlab.org/atlis/task/10652">ATLis
              10652</a>: for new arc optics from Arne.
            ORFP before/after installing new arc optics OWLs Tue-Wed Jan
            4-5.
            Some process bugs resolved, arc 1 looked pretty good, Arc 2
            did
            not in vertical. Backed out after FOPT. Arne has plan to
            refit,
            respecting Valeri's optics, in 2S triplet/quadruplet region.
            Arne doesn't touch extraction quads in any of these fits.
          </p>
        </div>
        <div class="outline-4">
          <h4 id="sec-25">2.2.5 Pathlength vs MOmod: Benesch</h4>
          <p>Studied during day 2010-01-05 during setup. NL response
            seems to
            be about 80% of SL. Haven't looked at owl MOmod. Looks like
            20 mV/deg
            in NL, 25 mV/deg in SL, different than last measurement.
            Will continue
            to evaluate.
          </p>
        </div>
      </div>
      <div class="outline-3">
        <h3 id="sec-26">2.3 Beam Studies Planning</h3>
        <div class="outline-4">
          <h4 id="sec-27">2.3.1 Mott Measurements: Grames/Suleiman (~2h)</h4>
          <p><a href="http://cebaf.jlab.org/atlis/task/10588">ATLis
              10589</a>: Complete characterization of new detectors
            after
            shutdown installation, spin rotator checkout.
          </p>
        </div>
        <div class="outline-4">
          <h4 id="sec-28">2.3.2 Qweak Injector Studies/Parity Quality
            Check: Mark Dalton</h4>
          <p><a href="https://cebaf.jlab.org/atlis/task/10625">ATLis
              10625</a>: Shifted to Thursday owl 2010-01-06 at Mark's
            request.
            But also should be revisited at high current.
          </p>
        </div>
        <div class="outline-4">
          <h4 id="sec-29">2.3.3 Prebuncher control module swap, injector
            retune</h4>
          <p><a href="https://cebaf.jlab.org/atlis/task/10508">ATLis
              10508</a>: To reduce 9.5 kHz noise when running prebuncher
            above 2. Was planned for 12/16. Not done; deferred to when?
            Not good confidence in spare; looks different physically.
            This ATLis does not seem strongly motivated until we go back
            to high injector voltage.
          </p>
          <ul>
            <li><span class="todo">TODO</span> Should investigate now
              we're at 100 kV w/lower prebuncher setting.<br>
            </li>
          </ul>
        </div>
        <div class="outline-4">
          <h4 id="sec-30">2.3.4 Determine multiharp masking requirements
            in tune/CW: Benesch<br>
          </h4>
          <p><a href="https://cebaf.jlab.org/atlis/task/10647">ATLis
              10647</a>: New ATLis, not scheduled yet. Mainly hall C.
          </p>
        </div>
        <div class="outline-4">
          <h4 id="sec-31">2.3.5 Tuesday 2011-01-11: (Test plans to be
            prepped)</h4>
          <p>Reza: 2-3h high current setup
            Tomasz Plawski: pre-buncher hardware calibration ~1-2h
            Alicia: Backphasing bunch length measurement @100kV, ~2h
            MikeS: Targeting 22nd for kicker commissioning, ready for
            beam afterwards
            MikeS: Have ~2h of work to do in Arc1, will prep test plan
          </p>
        </div>
      </div>
      <div class="outline-3">
        <h3 id="sec-32">2.4 Facility rad surveys, activation before
          shutdown: Yan</h3>
        <div class="outline-4">
          <h4 id="sec-33">2.4.1 <span class="todo">TODO</span> Discuss
            rad surveys, hot spots from just before shutdown.</h4>
          <p>NL rad survey 12/23 shows some elevated radiation areas,
            e.g. 1L26/27/28
            Survey 12/22 shows everything clean but end of NL marked as
            "hot spot"
            SL rad survey on 12/23 also shows some activation between
            22-23.
            This was when we went from 130 kV to 100 kV; after that
            switch beam
            archiver shows quite ugly; beam to hall A and Yves's test
            plan?
            FOPT on 12/22 1800 clean. Elog shows vacuum getting
            progressively
            worse from 12/22-23 – linac steering correction in 8R
            helped. Match
            probably needed redo from injector into NL after 130->100
            kV change.
            FOPT probably wouldn't have shown injector mismatch. This
            seems to
            imply that changing injector voltage requires a rematch. Use
            relative
            BLM screens and compare before/after.
          </p>
        </div>
      </div>
      <div class="outline-3">
        <h3 id="sec-34">2.5 Arc Diagnostics: Benesch</h3>
        <div class="outline-4">
          <h4 id="sec-35">2.5.1 <span class="todo">TODO</span> Analyze
            raw harp profile data and write up recommendations.</h4>
        </div>
      </div>
      <div class="outline-3">
        <h3 id="sec-36">2.6 12GeV Pathlength Requirements Status:
          Tiefenback</h3>
        <div class="outline-4">
          <h4 id="sec-37">2.6.1 <span class="todo">TODO</span> Need
            conclusions document.</h4>
        </div>
      </div>
      <div class="outline-3">
        <h3 id="sec-38">2.7 AOB?</h3>
        <div class="outline-4">
          <h4 id="sec-39">2.7.1 <span class="todo">TODO</span> Need
            volunteers for Sat 2011-01-22 USPAS afternoon tours</h4>
          <p>Michele and Jay have volunteered (2010-12-22 meeting); CASA
            folks
            have been surveyed for volunteers.
          </p>
        </div>
      </div>
    </div>
    <div id="postamble">
      <p class="author"> Author: Todd Satogata
        <a href="mailto:satogata@Todd-Satogatas-MacBook-Pro.local"><satogata@Todd-Satogatas-MacBook-Pro.local></a>
      </p>
      <p class="date"> Date: 2011/01/12 12:50:38</p>
    </div>
  </body>
</html>